OpenStudy (anonymous):

Calculate the average rate of change for the graphed sequence from n = 1 to n = 3.

OpenStudy (anonymous):

1 2 3 6

OpenStudy (amistre64):

the average rate of change is the slope of the line between end points ...

OpenStudy (amistre64):

the amount of y covered, in respect to the amount of x covered ....

OpenStudy (amistre64):

i dont see that as an option tho ... so you might need to provide more context

OpenStudy (amistre64):

duh ... 1 to 3; not 1 to 6 :)

OpenStudy (amistre64):

what is the slope from 1,2 to 3,8

OpenStudy (anonymous):

Right, so do I just use the slope formula for this?

OpenStudy (anonymous):

I wasnt sure because the y values arent very clear @amistre64

OpenStudy (amistre64):

slope formula is fine ... and yeah tiny pics on these old eyes kinda bite lol

OpenStudy (anonymous):

Okay, Im not getting any of the answers? (1,2) (3,8) 3-1/8-2=2/6=1/3 @amistre64

OpenStudy (anonymous):

Oops never mind, Y is supposed to be on top so the answer should be 3

OpenStudy (amistre64):

:) perfect